ICM 1998, Berlin, Aug. 18–27 Abstracts of Plenary and Invited Lectures

Section:

11. Mathematical Physics

1991 MS Classification: 81T30 Paul S. Aspinwall, Duke University, Durham, USA Geometry and String Duality (See section 4)

ICM 1998 Berlin · Plenary and Invited Lectures